OpenStudy (anonymous):

can someone help me with 4 questions A bag contains 4 white, 3 blue, and 5 red marbles Find the probability of choosing a red marble, then a white marble if the marbles are replaced.

OpenStudy (anonymous):

A 1/12 B 5/36 C 5/6 D 5/12

OpenStudy (anonymous):

@triciaal @welshfella @hatnn

OpenStudy (anonymous):

@SamsungFanBoy

OpenStudy (anonymous):

\[\frac{ red-marbles }{total-marbles }*\frac{ white-marbles }{ total-marbles }\]

OpenStudy (anonymous):

what????

OpenStudy (anonymous):

How many red marbles are there?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

How many total marbles?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

So red-marbles/total-marbles would be what?

OpenStudy (anonymous):

5/12

OpenStudy (anonymous):

Now do the same thing with white marbles and total marbles.

OpenStudy (anonymous):

4/12

OpenStudy (anonymous):

which would simplify to?

OpenStudy (anonymous):

5/6

OpenStudy (anonymous):

How does 4/12 simplify to 5/6?

OpenStudy (anonymous):

sorry I mean 1/12

OpenStudy (anonymous):

How does 4/12 simplify to 1/12?

OpenStudy (anonymous):

A 1/12 B 5/36 C 5/6 D 5/12 A 1/12 B 5/36 C 5/6 D 5/12 A 1/12 B 5/36 C 5/6 D 5/12

OpenStudy (anonymous):

Uh, what?

OpenStudy (anonymous):

these are the chocies to chose from

OpenStudy (anonymous):

Ok, but let's get back to what we're working on. What does 4/12 simplify to?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

What does 4/4 equal? What does 12/4 equal?

OpenStudy (anonymous):

4/4 = 1

OpenStudy (anonymous):

And 12/4?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

So \[\frac{ 4/4 }{ 12/4 }\] would be what?

OpenStudy (anonymous):

but 1/3 is not a chioic to chose from

OpenStudy (anonymous):

A 1/12 B 5/36 C 5/6 D 5/12

OpenStudy (anonymous):

Ok, now we have the fractions. 5/12 and 1/3. Now multiply the fractions.

OpenStudy (anonymous):

my calculator doesn't do fractions

OpenStudy (anonymous):

You don't need a calculator for this. It's just\[\frac{ 5 }{ 12 }*\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

5/36

OpenStudy (anonymous):

And that is your answer.

OpenStudy (anonymous):

so the answer is b 5/36 right

OpenStudy (anonymous):

Yes, I just said so.

OpenStudy (anonymous):

can you help me with the other 3 plz

OpenStudy (anonymous):

I will if you tag me in a separate question.

OpenStudy (anonymous):

using the same info A bag contains 4 white, 3 blue, and 5 red marbles Find the probability of choosing 3 blue marbles in a row if the marbles are replaced

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

hello u still there

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@Loser66 and legendarsadist will u help me with the three questions plz plz

OpenStudy (anonymous):

using the same info A bag contains 4 white, 3 blue, and 5 red marbles Find the probability of choosing 3 blue marbles in a row if the marbles are replaced

OpenStudy (loser66):

The probability of blue marbles is 1/4 for each time, right? (That is from \(\dfrac{blue}{total}\)

OpenStudy (anonymous):

A 2/11 B 1/220 C 1/27 D 1/64

OpenStudy (loser66):

After pick it out, then replace, second time around, the situation repeats. This time, the result is exactly what it happen in the first time: 1/4

OpenStudy (loser66):

Third time, same

OpenStudy (loser66):

Hence, the final result is (1/4)*(1/4)*(1/4) = ????

OpenStudy (anonymous):

1/64

OpenStudy (loser66):

yup

OpenStudy (anonymous):

so the answer is d

OpenStudy (loser66):

yup